# Soft Reasoning on Uncertain Knowledge Graphs

### v1

@@ Appendix B Constraint Satisfaction Problem and Soft Constraint Satisfaction Problems
Many problems can be viewed as CSP, which include workforce scheduling and the toy 8-queens problem. Conjunctive queries are a special case to be reduced as CSP under open world assumptions if we set the constant variable’s domain as itself, set the domain of the existential variable and free variable as the entity set $\mathcal{E}$ of knowledge graphs, and treat atomic formula or its negation as binary constraint by knowledge graph.
@@
Though CSP is a very powerful formulation, it fails when real-life problems need to describe the preference of constraint. It usually returns null answers for problems with many constraints, which are called over-constraints. To tackle the above weakness, many versions of Soft Constraint Satisfaction Problems (SCSP) are developed, such as fuzzy CSP, weighted CSP, and probabilistic CSP, which all follow a common semi-ring structure, where two semi-ring operations ×s,+$\times_{s},+_{s}$ are utilized to model constraint projection and combination respectively. Based on this theoretical background, we propose soft queries based on SCSP. The proposed soft queries will have the advantages of SCSPs and can handle the numeric facts representing uncertainty.
Appendix C Details of Implementation
C.1 Uncertain knowledge graph embedding
We reproduce the results of Chen et al. (2019, 2021a) and use the same embedding dimension. We search the other parameters including the learning rate from $\{1e-3,5e-4,1e-4\}$ and regularization term $\lambda$ from $\{0.1,0.01,0.05\}$.
C.2 Query embedding with number embedding
@@ C.3 Two strategies of calibration
Appendix D Details of Soft Reasoning with Calibrated Confidence Values
@@
D.1 Uncertain value definition
@@ D.2 Details of the inference of SRC
D.2.2 Step 2. Remove constant nodes.
@@
D.2.3 Step 3. Remove self-loop edges.
@@
D.2.4 Step 4. Remove leaf nodes.
@@
D.2.5 Step 5. Enumerate on the cycle.
@@
D.2.6 Step 6. Getting the utility vector.
@@
Appendix E Uncertain Knowledge Graph Embeddings
@@
UKGE (Chen, 2023) is a vector embedding model designed for uncertain knowledge graphs. It has been tested on three tasks: confidence prediction, relational fact ranking, and relational fact classification. To address the sparsity issue in the graph, UKGE utilizes probabilistic soft logic, allowing for the inclusion of additional unseen relational facts during training.
BEUrRE (Chen et al., 2021a) is a probabilistic box embedding model that has been evaluated on two tasks: confidence prediction and relational fact ranking. This model represents each entity as a box and captures relations between two entities through affine transformations applied to the head and tail entity boxes.
@@
Appendix F Float Embedding for query Embedding
@@ F.1 Float embedding
F.2 Modified relation projection
@@
Appendix G Details in the Main Dataset Construction
G.1 Uncertain knowledge graphs
Table 7: The Statistics of the knowledge graphs. We provided the counts of each knowledge graph’s nodes and edges. Additionally, we presented the mean and standard deviation of the scores.
Uncertain KG Entities Relations Facts AVG. STD.
CN15k 15,000 36 204984 0.629 0.231
PPI5k 4999 7 230929 0.415 0.213
O*NET20k 20,643 19 418304 0.301 0.260
@@ G.1 Uncertain knowledge graphs
O*NET20K (Pai & Costabello, 2021) is a subset of O*NET, a dataset that describes labeled binary relations between job descriptions and skills. The associated values are to evaluate the importance of the link within the triple.
We present the statistics of these three uncertain knowledge graphs in Table 7.
G.2 Soft requirements
@@ G.4 Evaluation protocol
Appendix H Details in Varying Soft Requirements Setting
In this setting, we aim to test the model’s generalization ability on soft requirements under different strategies. We have chosen a pair of different strategies for each of the two parameters, resulting in a total of four groups. Specifically, we selected ”zero” and ”normal” for parameter ”a,” and ”equal” and ”normal” for parameter ”b.” Detailed statistics for this setting can be found in the table below. For each group, we follow the procedure and sample train, valid, and test queries. We present the additional results in Table 10.
@@ Appendix H Details in Varying Soft Requirements Setting
Appendix I Details of Large Language Model Evaluation Setting
@@
Table 11: The number of annotated queries.
I.1 Prompt
{spverbatim}
# Background Given a soft query containing a free variable f, there are four candidate entities for this variable and you need to choose the best of them to satisfy the soft query most. In order to do so, you can first compute the
confidence value to see whether the chosen candidate entity satisfies the soft query after substituting the
free variable with a given candidate entity. After computing the confidence values for each corresponding candidate entity, you need to pick the entity that leads to the highest confidence value. The detailed steps are described below:
## Definition of Soft Atomic Constraint: A soft atomic constraint c, a.k.a. a soft atomic formula or its negation, is in the form of (h, r, t, α, β) or
neg (h, r, t, α, β).
### Notation Description: In each constraint, we have four different types of variables. 1. r is a relation; 2. h and t are two terms. Each term represents an entity or a variable whose values are entities. And free variable is a term. 3. αis called the necessity value, which represents the minimal requirement of the uncertainty degree of this constraint. It can be any decimal between 0.0 and 1.0. If the confidence value is less than the necessity value, the constraint is not satisfied, and thus the final confidence value becomes negative infinity; 4. βrepresents the priority of this constraint and can be any decimal between 0.0 and 1.0.
### Confidence Value of Soft Constraints V(c): 1. The triple (h,r,t) comes from the relation fact in the knowledge graph. In our setting, the relation fact r(h,t) is not boolean, and it has a confidence value in the range [0.0,1.0] according to its plausibility. When the constraint is negative, you need to first estimate r(h,t)—the confidence value of r(h,t)—and use 1-r(h,t) as the final confidence value for this negative constraint. 2. αis the threshold in our filter function, working as follows: f(v, α) = v, if v ≥α, -∞, if v ¡ α. 3. βis a coefficient.
Thus, the final equation becomes: V(h, r, t, α, β) = β×f(r(h,t), α), V(
neg (h, r, t, α, β)) = β×f(1-r(h,t), α).
## Definition of Conjunctive Queries ϕ: Soft conjunctive queries are composed of soft constraints.
### Notation Description: Conjunctive query ϕ= c_1 ∧…∧c_n, where c_i is a soft constraint.
### Confidence Value of Soft Conjunctive Queries V(ϕ) : First, you need to compute the confidence values of all soft constraints in this soft conjunctive query. Then, you can simply sum up these confidence values as the final confidence value of the soft conjunctive query as follows: V(ϕ) = ∑_i V(c_i).
If the conjunctive query has an existential variable e, you should find an entity to replace it and then compute the confidence value of this query.
# Output Format Please output your response in the JSON format, where the first element is the best candidate entity among the four options, and the second element is your explanation for your choice.
# Question Soft query:(h_1,r_1,f,α_1,β_1) ∧(h_2,r_2,f,α_2,β_2) Four candidate entities: s_1, s_2, s_3, s_4
@@ I.1 Prompt
Appendix J Proof for Error Analysis
J.1 Proof of Theorem 1
@@
$\Pr\left(\|\hat{U}[\psi](s)-\mathcal{U}[\psi](s)\|>\delta)\right)=\Pr\left(\beta\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\delta\right)=\Pr\left(\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\frac{\delta}{\beta}\right)$
@@
$\Pr\left(\|\hat{U}[\psi](s)-\mathcal{U}[\psi](s)\|>\delta)\right)=\Pr\left(\beta\|[1-\hat{P}(s,r,o)]_{\alpha}+[\mathcal{P}(s,r,o)]_{\alpha}-1\|>\delta\right)=\Pr\left(\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\frac{\delta}{\beta}\right)$

### v2

@@ Appendix F Constraint Satisfaction Problem and Soft Constraint Satisfaction Problems
Many problems can be viewed as CSP, which include workforce scheduling and the toy 8-queens problem. Conjunctive queries are a special case to be reduced as CSP under open-world assumptions if we set the constant variable’s domain as itself, set the domain of the existential variable and free variable as the entity set $\mathcal{E}$ of knowledge graphs, and treat atomic formula or its negation as binary constraint by knowledge graph.
@@
Though CSP is a very powerful formulation, it fails when real-life problems need to describe the preference of constraint. It usually returns null answers for problems with many constraints, which are called over-constraints. To tackle the above weakness, many versions of Soft Constraint Satisfaction Problems (SCSP) are developed, such as fuzzy CSP, weighted CSP, and probabilistic CSP, which all follow a common semiring structure, where two semiring operations ×s,+$\times_{s},+_{s}$ are utilized to model constraint projection and combination respectively. Based on this theoretical background, we propose soft queries based on SCSP. The proposed soft queries will have the advantages of SCSPs and can handle the numeric facts representing uncertainty.
Appendix G Details of Implementation
Our experiments are run on the Nvidia V100-32G.
G.1 adaptive scoring
Let $E_{s}$, $E_{r}$, and $E_{t}$ be the embedding vectors of entity $s$, relation $r$, and entity $t$ respectively. We parameterize the adaptive scoring calibration using the following learnable affine transformations:
$\rho_{\theta}(s,r,o)$ $=W^{1}_{1}E_{s}+b^{1}_{1}+W^{1}_{2}E_{r}+b^{1}_{2}+W^{1}_{3}E_{t}+b^{1}_{3},$ (48)
$\lambda_{\theta}(s,r,o)$ $=W^{2}_{1}E_{s}+b^{2}_{1}+W^{2}_{2}E_{r}+b^{2}_{2}+W^{2}_{3}E_{t}+b^{2}_{3},$ (49)
where $\{W^{i}_{j},b^{i}_{j}\}$ for $1\leq i\leq 2$ and $1\leq j\leq 3$ are the learnable parameters.
G.2 Uncertain knowledge graph embedding
We reproduce previous results (Chen et al., 2019, 2021a) and use the same embedding dimension. We search the other parameters including the learning rate from $\{1e-3,5e-4,1e-4\}$ and regularization term $\lambda$ from $\{0.1,0.01,0.05\}$.
G.3 Query embedding with number embedding
@@ G.4 Two strategies of calibration
Appendix H Details of Soft Reasoning with Calibrated Confidence Values
@@
H.1 Uncertain value definition
@@ H.2 Details of the inference of SRC
H.2.2 Step 2. Remove constant nodes.
@@
H.2.3 Step 3. Remove self-loop edges.
@@
H.2.4 Step 4. Remove leaf nodes.
@@
H.2.5 Step 5. Enumerate on the cycle.
@@
H.2.6 Step 6. Getting the utility vector.
@@
Appendix I Uncertain Knowledge Graph Embeddings
@@
UKGE Chen (2023) is a vector embedding model designed for uncertain knowledge graphs. It has been tested on three tasks: confidence prediction, relational fact ranking, and relational fact classification. To address the sparsity issue in the graph, UKGE utilizes probabilistic soft logic, allowing for the inclusion of additional unseen relational facts during training.
BEUrRE Chen et al. (2021a) is a probabilistic box embedding model that has been evaluated on two tasks: confidence prediction and relational fact ranking. This model represents each entity as a box and captures relations between two entities through affine transformations applied to the head and tail entity boxes.
@@
Appendix J Float Embedding for query Embedding
@@
J.2 Modified relation projection
@@
Appendix K Details in the Main Dataset Construction
Figure 4: Query structures of query types. The white, yellow, and red circles represent constant, existential, and free nodes, respectively. The negative atomic formulas are represented by red edges, while atomic formulas are represented by black edges. Like the previous naming convention (Ren and Leskovec, 2020; Yin et al., 2024), we use ”P” for projection, “I” for intersection, “N” for negation, “M” for multi-edge, and “L” for existential leaf.
K.1 Uncertain knowledge graphs
@@
O*NET20K (Pai and Costabello, 2021) is a subset of O*NET, a dataset that describes labeled binary relations between job descriptions and skills. The associated values are to evaluate the importance of the link within the triple.
We present the statistics of these three uncertain knowledge graphs in Table LABEL:tab:_Statistics_of_graph.
K.2 Soft requirements
@@
Appendix L Details in Varying Soft Requirements Setting
In this setting, we aim to test the model’s generalization ability on soft requirements under different strategies. We have chosen a pair of different strategies for each of the two parameters, resulting in a total of four groups. Specifically, we selected "zero" and "normal" for parameter "a," and "equal" and "normal" for parameter "b." Detailed statistics for this setting can be found in the table below. For each group, we follow the procedure and sample train, valid, and test queries. We present the additional results in Table 10.
@@
Appendix M Details of Large Language Model Evaluation Setting
@@
M.1 Prompt
# Background
Given a soft query containing a free variable f, there are four candidate entities for
this variable and you need to choose the best of them to satisfy the soft query most. In
order to do so, you can first compute the confidence value to see whether the chosen
candidate entity satisfies the soft query after substituting the free variable with a
given candidate entity. After computing the confidence values for each corresponding
candidate entity, you need to pick the entity that leads to the highest confidence value.
The detailed steps are described below:
## Definition of Soft Atomic Constraint:
A soft atomic constraint c, a.k.a. a soft atomic formula or its negation, is in the form
[…]
### Notation Description:
In each constraint, we have four different types of variables.
1. r is a relation;
2. h and t are two terms. Each term represents an entity or a variable whose values
are entities. And free variable is a term.
[…]
of the uncertainty degree of this constraint. It can be any decimal between 0.0 and
1.0. If the confidence value is less than the necessity value, the constraint is
not satisfied, and thus the final confidence value becomes negative infinity;
[…]
between 0.0 and 1.0.
### Confidence Value of Soft Constraints V(c):
1. The triple (h,r,t) comes from the relation fact in the knowledge graph. In our
setting, the relation fact r(h,t) is not boolean, and it has a confidence value
in the range [0.0,1.0] according to its plausibility. When the constraint is
negative, you need to first estimate r(h,t)the confidence value of r(h,t)and
use 1-r(h,t) as the final confidence value for this negative constraint.
[…]
Thus, the final equation becomes:
[…]
Soft conjunctive queries are composed of soft constraints.
### Notation Description:
[…]
First, you need to compute the confidence values of all soft constraints in
this soft conjunctive query. Then, you can simply sum up these confidence
values as the final confidence value of the soft conjunctive query as follows:
[…]
If the conjunctive query has an existential variable e, you should find an
entity to replace it and then compute the confidence value of this query.
# Output Format
Please output your response in the JSON format, where the first element
is the best candidate entity among the four options, and the second element
is your explanation for your choice.
# Question
[…]
Four candidate entities: s_1, s_2, s_3, s_4
@@ 1.0. If the confidence value is less than the necessity value, the constraint is
Appendix N Proof for Error Analysis
N.1 Proof of Theorem 1
@@
$\Pr\left(\|\hat{U}[\psi](s)-\mathcal{U}[\psi](s)\|>\delta)\right)$ $=\Pr\left(\beta\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\delta\right)$ (67)
$=\Pr\left(\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\frac{\delta}{\beta}\right)$ (68)
@@
$\Pr\left(\|\hat{U}[\psi](s)-\mathcal{U}[\psi](s)\|>\delta)\right)$ $=\Pr\left(\beta\|[1-\hat{P}(s,r,o)]_{\alpha}+[\mathcal{P}(s,r,o)]_{\alpha}-1\|>\delta\right)$
$=\Pr\left(\|[\hat{P}(s,r,o)]_{\alpha}-[\mathcal{P}(s,r,o)]_{\alpha}\|>\frac{\delta}{\beta}\right)$